# Optimal Sizing and Material Choice for Additively Manufactured Compact Plate Heat Exchangers

### v1

@@ Abstract
With advancements in additive manufacturing (AM) capabilities, new opportunities arise to design compact heat exchangers (cHEXs) that leverage AM’s degrees of freedom (DOFs) to enhance energy and material efficiency. However, excessive size reduction in counterflow cHEXs can compromise effectiveness due to axial heat conduction through the solid material, which is strongly influenced by thermal conductivity and wall thickness. This study investigates how AM material selection and related limitations in AM thin-wall production might compromise the core size of counterflow plate heat exchangers when targeting maximum power density. An optimization framework evaluates maximum power densities for six materials, i.e., plastic, austenitic steel, Al2O3, AlN, aluminum, and copper. The evaluations are conducted under constant effectiveness and pressure drop while accounting for AM-specific plate thickness limitations and a lower bound on plate spacing to address fouling concerns. Across all scenarios, copper cHEXs exhibit the lowest maximum power density, despite their high thermal conductivity. Without constraints on plate thickness and spacing, the optimal plastic heat exchanger achieves a power density 1800 times greater than the steel baseline, while copper shows a decrease by a factor of 0.98. With an equal plate thickness of 0.5 mm for all materials, plastic retains the highest power density, achieving 12.2 times more than copper. Introducing a fouling constraint of 0.8 mm plate spacing shifts the optimal material to austenitic steel. When material-specific plate thicknesses are considered, the plastic cHEX achieves the highest power density, five times greater than copper, due to superior thin-wall resolution. This study highlights the significant impact of AM constraints on the energy and material efficiency of cHEXs. It also demonstrates that materials with lower thermal conductivity, such as plastic or austenitic steel, can outperform high-conductivity materials like copper in compact heat exchanger designs.
@@
Heat exchanger; counterflow; axial wall conduction; design optimization; additive manufacturing; manufacturing constraints.
@@ 1 Introduction
Heat exchangers are crucial components in various engineering fields, including the automobile and aerospace industries, microelectronics cooling, and energy conversion systems. Increasing demands for compact and lightweight solutions, particularly in mobile applications, have intensified the need for high-performance compact heat exchangers (cHEXs).
@@
Producing cHEXs using conventional manufacturing (CM) poses significant challenges. CM methods, such as subtractive techniques and forming processes, often require the assembly of components through welding or brazing, making it difficult to achieve leak-proof designs. Additionally, CM’s limitations in machining complex geometries further complicate the production of compact, high-performance heat exchangers [1].
Additive manufacturing (AM) has emerged as a transformative technology for cHEX production, offering unparalleled freedom in design complexity and the ability to manufacture components in a single run [2, 3, 4]. This capability reduces the risk of fluid leakage, minimizes material waste, and supports strength-critical designs [1, 5]. These outstanding features have gathered significant interest among heat exchanger designers, driving ongoing efforts to develop lightweight cHEX using AM technology.
@@
Despite AM’s promise, axial heat conduction remains a critical issue in cHEX design. As heat exchanger volume decreases, axial conduction within the solid plate separating the fluids becomes more significant, reducing thermal effectiveness [6, 7, 8, 5]. This phenomenon can dominate performance, potentially offsetting the benefits of AM-enabled design enhancements.
@@
While extensive research exists on AM cHEXs and axial conduction effects, these topics are often addressed separately. Studies on AM processes frequently prioritize high thermal conductivity materials, such as aluminum and copper [4, 9]. Conversely, research on axial wall conduction emphasizes that high thermal conductivity may negatively impact the thermal efficiency of cHEX [10, 11, 12]. Additionally, thin plates are known to mitigate axial conduction [13], but AM processes impose constraints on minimum achievable plate thickness compared to CM. A comprehensive evaluation of axial conduction in AM cHEXs, accounting for these limitations, remains unexplored.
Several studies have demonstrated the potential of AM in heat exchanger design. For instance, Saltzman et al.[14] compared conventionally manufactured plate-fin air-liquid crossflow heat exchangers with AM-produced counterparts, showing improvements in heat transfer performance but noting higher pressure drops due to AM surface roughness. Da Silva et al.[15] and Moon et al. [16] further demonstrated the feasibility of AM for producing compact and efficient heat exchangers, highlighting its ability to fabricate geometries unattainable with CM. However, these studies do not fully address the interplay between material properties, axial conduction, and AM constraints.
@@
For general (non-AM) cHEXs, axial conduction has been extensively studied theoretically [6, 17, 18, 19], numerically [20, 21, 22, 23], and experimentally [24, 25]. Researchers have examined its effects on thermal performance under various conditions, emphasizing the importance of parameters such as thermal conductivity, plate thickness, and plate spacing [13, 26]. Optimization studies, such as those by Rogiers and Baelmans [27] and Buckinx et al. [28], have further explored the effect of geometric parameters on the power density. These studies provide valuable insights but often assume ideal manufacturing conditions without considering AM-specific constraints.
In this paper, we investigate the influence of material thermal properties on the design of additively manufactured cHEX and evaluate the impact of axial wall conduction, considering the limitations of AM in producing thin walls. The goal is to achieve the minimum size of a heat exchanger without compromising on pressure drop and thermal effectiveness. When thermal power is fixed, achieving the minimum size corresponds to a maximal power density, which is thus defined as the design objective. The study simplifies the analysis by focusing on a parallel flat plate counterflow heat exchanger, using length and plate spacing as design variables and plate thickness as a design constraint. Unlike the typical heat exchanger design approach, where plate thickness and spacing are predetermined and sizing is adjusted to fit available space [29], this alternative approach explores how plate thickness, spacing, and length can be chosen to optimize power density. In the optimizations, the constant pressure drop and effectiveness are first determined based on a baseline heat exchanger. Then, the maximal power density geometry is calculated considering AM constraints for six different materials. To address the risk for fouling in maximal power density designs, a minimum plate spacing constraint (“fouling constraint”) is introduced and its impact is studied. Furthermore, to examine the effect of effectiveness and plate thickness on power density, optimizations are repeated under various effectiveness and plate thickness conditions. Hence, this comprehensive evaluation discusses the impact of AM constraints on the design of cHEX manufactured with different materials.
@@ 2 Thermal model
The present investigation is conducted on a simple parallel flat plate heat exchanger (PFPHE), as illustrated in Figure 1. The configuration comprises $12n+$ 1 flat, thin plates arranged in parallel, with each plate stacked on top of the preceding one. The plates have a thickness of $t$ and form $2n$ fluid channels in a repeating sequence, alternating between cold and hot flows. The channels are of identical spacing $D$, length $L$, and width $W$. While this type of heat exchanger can be configured with various flow arrangements, including parallel flow, crossflow, and counterflow, only counterflow is considered in this investigation. Furthermore, this study assumes a perfectly balanced heat exchanger with equal mass flow rates, identical pressure drops, and uniform fluid properties on both the hot and cold sides.
@@
The goal of this study is to minimize the size of a parallel-plate counterflow heat exchanger without compromising pressure drop and thermal effectiveness. For a given thermal power, this implies maximizing thermal power density. This section provides a comprehensive description of the thermal-hydraulic model used to evaluate power density taking axial heat conduction into account.
@@
where $\varepsilon$ is the effectiveness, representing the ratio of the actual heat transfer rate to the maximum possible heat transfer rate [30], $\dot{m}$ is the mass flow rate, $c_{p}$ is the fluid specific heat at constant pressure, and $\Delta T$ is the temperature difference between the hot and cold side inlets. Assuming a sufficiently large number of plates, the total height of the heat exchanger is approximated as $H=(2n)D+(2n+1)t\approx 2nD(1+t/D)$.
@@
where $f$ is the Fanning friction factor, Re is the Reynolds number, and $\mu$ is the dynamic viscosity. Upon substituting Eq. 3 into Eq 2, the mass flow rate $\dot{m}$ becomes
@@
Subsequently, by substituting Eq. 4 into Eq. 1 and performing some rearrangements, an analytical relation for a non-dimensional power density $Q^{\diamond}$ is obtained:
$Q^{\diamond}\equiv\frac{\mathcal{D}\mu}{c_{p}\Delta T\rho\Delta P_{\text{ref}}}=\psi\frac{1}{f\textup{Re}}\frac{1}{1+t/D}\varepsilon\left(\frac{D}{L}\right)^{2},$ (5)
where $\psi$ represents the ratio of the actual pressure drop $\Delta P$ to a reference pressure drop $\Delta P_{\text{ref}}$. Throughout this work $\Delta P_{\text{ref}}$ is set equal to the pressure drop of the initial design given in Table 1.
Kroeger [6] derived an expression to calculate the effectiveness of a balanced counterflow heat exchanger by taking the axial heat conduction in the wall into account.
@@
$\phi=\sqrt{\frac{\textup{M}\cdot\textup{NTU}}{1+\textup{M}\cdot\textup{NTU}}}\tanh{\frac{\textup{NTU}}{\sqrt{\textup{M}\cdot\textup{NTU}/(1+\textup{M}\cdot\textup{NTU}}}}.$ (7)
The number of transfer units NTU is a dimensionless parameter that combines the dimensional terms of the overall heat transfer coefficient $U$, heat transfer surface area $A$, mass flow rate $\dot{m}$, and specific heat $c_{p}$ into a single term [31], i.e. $\textup{NTU}=\frac{UA}{\dot{m}c_{p}}$. Meanwhile, the axial conduction parameter $\textup{M}=\frac{UA_{ax}}{\dot{m}c_{p}}$ addresses the effect of heat conduction along the heat exchanger wall [6, 20], with $A_{ax}$ is the heat transfer surface area in the axial direction.
@@
where $h$ is the convective heat transfer coefficient, and $k_{w}$ is the thermal conductivity of the plate. The first term within the brackets accounts for the combined convective resistances of the hot and cold flow sides, while the second term addresses the lateral conduction resistance through the plate. By substituting the expression for the convective heat transfer coefficient $h=\frac{\textup{Nu}k}{2D}$, $UA$ is transformed into
$UA=2nk\frac{L}{D}\frac{\textup{Nu}}{4}W\left(1+\frac{\textup{Nu}}{4}\frac{k}{k_{w}}\frac{t}{D}\right)^{-1},$ (9)
with Nu is the Nusselt number. The axial plate conductance per unit length $UA_{ax}$, which is used in the calculation of the axial conduction parameter M, is defined as
@@
Rogiers and Baelmans [27] formulated the number of transfer units NTU and the axial conduction parameter M as functions of the dimensionless Bejan number $\textup{Be}=\frac{\Delta PL^{2}}{\mu\alpha}$ and relative plate spacing $\widetilde{D}=\frac{D}{L}$. To this end, Eqs. 4 and 9 are substituted into the NTU definition, while Eqs. 4 and 10 are substituted into the M definition:
$\textup{NTU}(\widetilde{D},\textup{Be})=\frac{f\textup{Re}\textup{Nu}}{4}\left(1+\frac{\textup{Nu}}{4}\frac{k}{k_{w}}\frac{t}{D}\right)^{-1}\widetilde{D}^{-4}\textup{Be}^{-1},$ (11)
$\textup{M}(\widetilde{D},\textup{Be})=f\textup{Re}\frac{k_{w}}{k}\frac{t}{D}\widetilde{D}^{-2}\textup{Be}^{-1}.$ (12)
As a result, by combining Eqs. 11, 12, and Eq. 6, the effectiveness $\varepsilon$ can be analytically expressed as a function of the Bejan number Be, which can be interpreted as the dimensionless plate length, and $\widetilde{D}$ as the dimensionless plate spacing.
Within the scope of this paper, only heat exchangers operating in the fully developed laminar flow regime are investigated. For a laminar channel flow between parallel planes, $f\textup{Re}$=24 [32]. The flow is assumed to be a two-dimensional (2D) planar channel flow, and it is also assumed that the Nusselt number Nu is known prior to optimization. For a 2D planar channel flow subjected to constant surface heat flux, Nu=8.235 [32].

### v2

@@ Abstract
Advances in additive manufacturing (AM) enable new opportunities to design compact heat exchangers (cHEXs) by leveraging flexible geometries to improve energy and material efficiency. However, it is well-known that reducing size in counterflow cHEXs can degrade effectiveness due to axial heat conduction through the solid material, which depends strongly on material thermal conductivity and wall thickness. Understanding this interaction between fundamental heat transfer mechanisms and manufacturing constraints is essential for designing next-generation compact thermal systems with the aspiration of taking full advantage of AM’s shaping flexibility. This study investigates how material selection and AM thin-wall limitations influence the maximum achievable power density in compact plate heat exchangers. An optimization framework evaluates six materials: plastic, austenitic steel, Al2O3, AlN, aluminum, and copper. The evaluations are performed under fixed pressure drop and effectiveness, while accounting for AM-specific thickness constraints and a lower limit on plate spacing to address fouling risks. Results show that copper consistently yields the lowest power density, despite having the highest thermal conductivity of the materials considered. In contrast, plastic achieves the highest power density across the large majority of optimization scenarios. Without manufacturing or fouling constraints, plastic outperforms the baseline steel design by a factor of 962. With a uniform plate thickness of 0.5 mm, it remains the top performer with 21.6 times the power density of copper. Adding a fouling constraint narrows the performance gap, making austenitic steel and plastic comparable. When material-specific thicknesses are applied, plastic again leads in compactness with 7.5 times the power density of copper, due to its superior thin-wall manufacturability. These findings highlight that AM constraints substantially affect cHEX compactness, and that lower-conductivity materials like plastic and austenitic steel can outperform metals like copper in power-dense heat exchanger designs. The results demonstrate how manufacturing limits, material properties, and heat transfer fundamentals jointly determine achievable compactness in AM-based heat exchangers.
@@
Compact plate heat exchanger; axial heat conduction; design optimization; additive manufacturing; manufacturing constraints.
Nomenclature
  • AA
    heat transfer surface area [m2]
  • $A_{ax}$
    heat transfer surface area in the axial direction [m2]
  • $c_{p}$
[…]
  • $D$
    plate spacing [m]
  • $\mathcal{D}$
    thermal power density, $\dot{q}/\left(HLW\right)$ [W/m3]
  • $D^{\ast}$
    nondimensional plate spacing, $D/t_{\text{ref}}$ [-]
  • $f$
    Fanning friction factor [-]
  • $H$
    total heat exchanger height [m]
  • IF
    Improvement Factor [-]
  • $k$
[…]
  • $k_{w}$
[…]
  • $L$
    plate length, channel length [m]
  • $L^{\ast}$
    nondimensional plate length, $L/t_{\text{ref}}$ [-]
  • $\dot{m}$
    mass flow rate at each side [kg/s]
  • M
    axial conduction parameter, $UA_{ax}/\left(\dot{m}c_{p}\right)$ [-]
  • $\dot{m}^{\prime}_{\text{single}}$
[…]
  • $n$
    number of hot side or cold side channels [-]
  • NTU
    number of transfer units, $UA/\left(\dot{m}c_{p}\right)$ [-]
  • Nu
    mean Nusselt number, h2D/k [-]
  • $\Delta P$
    pressure drop between inlet and outlet [Pa]
  • $\dot{q}$
    heat transfer rate [W]
  • $Q^{\diamond}$
    nondimensional power density, Eq 5 [-]
  • Re
    Reynolds number based on 2D, $\rho V2D/\mu$ [-]
  • $t$
    plate thickness [m]
  • $t^{\ast}$
    nondimensional plate thickness, $t/t_{\text{ref}}$ [-]
  • $\Delta T$
    hot to cold side inlet temperature difference [K]
  • $U$
[…]
  • $V$
    fluid bulk velocity in parallel-plate channel [m/s]
  • $W$
    plate width, channel width [m]
  • Greek symbols
  • $\alpha$
    fluid thermal diffusivity, $k/\left(\rho c_{p}\right)$ [m2/s]
  • $\varepsilon$
    heat exchanger effectiveness [-]
  • $\gamma$
    plate thickness-to-spacing ratio constant, $t^{\ast}/D^{\ast}$ [-]
  • $\mu$
[…]
  • $\Pi$
    dimensionless scaling group $\Delta P_{\text{ref}}t_{\text{ref}}^{2}/\left(\alpha\mu\right)$ [-]
  • $\psi$
    ratio of the actual pressure drop to the reference value, $\Delta P/\Delta P_{\text{ref}}$ [-]
  • $\rho$
    fluid density [kg/m3]
  • Subscripts
  • d
    design value
  • min
    minimum allowable value
  • ref
    reference value
@@ 1 Introduction
Heat exchangers are crucial components in various engineering fields, including the automotive and aerospace industries, microelectronics cooling, and energy conversion systems. Increasing demands for compact and lightweight solutions, particularly in mobile and decentralized energy applications, have amplified the need for high-performance compact heat exchangers (cHEXs) Wang2024 , Broatch2024 . As heat transfer applications move toward higher power density and smaller scales, connecting fundamental heat transfer behavior with practical design and manufacturing methods becomes increasingly important.
@@
Producing cHEXs using conventional manufacturing (CM) poses significant challenges. CM methods, such as subtractive techniques and forming processes, often require the assembly of components through welding or brazing, making it difficult to achieve leak-proof designs. Additionally, CM’s limitations in machining complex geometries further complicate the production of compact, high-performance heat exchangers Kaur2021 .
Additive manufacturing (AM) has emerged as a transformative technology for cHEX production, offering unparalleled freedom in design complexity and the ability to manufacture components in a single run Jafari2018 , Chen2021 , Careri2023 . This capability reduces the risk of fluid leakage, minimizes material waste, and supports strength-critical designs Kaur2021 , Fuchs2022 . Several studies have demonstrated the feasibility of AM for producing compact and efficient heat exchangers Saltzman2018 , highlighting its ability to fabricate geometries unattainable with CM DaSilva2021 , Moon2021 , Rao2025 , Choudhary2025 . Recent investigations have also employed topology optimization to improve heat transfer in additively manufactured or miniaturized thermal systems Marshall2022 , Alshayji2025 , underscoring the convergence of heat-transfer fundamentals and manufacturing innovation. Such approaches show how advanced design tools and AM can be combined to translate thermal optimization concepts into practical engineering solutions.
@@
However, despite these advantages, axial heat conduction remains a critical issue in cHEX design. As heat exchanger volume decreases, axial conduction within the solid wall separating the fluids becomes more significant, reducing thermal effectiveness Kroeger1967 , Nagasaki2003 , Maranzana2004 , Fuchs2022 and potentially offsetting the benefits of AM-enabled design enhancements. A better understanding of this effect requires considering both the underlying heat transfer mechanisms and the limitations imposed by AM processes.
@@
While extensive research exists on AM cHEXs and on axial conduction effects, these topics are often addressed independently. Studies on AM processes frequently prioritize high-conductivity materials, such as aluminum and copper Careri2023 , Hein2021 . Conversely, research on axial wall conduction emphasizes that high thermal conductivity may negatively impact the thermal efficiency of cHEXs Bier1993 , Stief1999 , Borjigin2018 . Additionally, thin plates are known to mitigate axial conduction Mitra2023 , but AM processes impose constraints on minimum achievable plate thickness due to limitations in manufacturing resolution, that is, the smallest reliably producible wall thickness specific to each AM technique and material in order to guarantee the cHEX to be leakage-tight. A comprehensive evaluation of axial conduction in AM cHEXs, accounting for these practical limitations, remains unexplored.
@@
For general (non-AM) cHEXs, axial conduction has been extensively studied theoretically Kroeger1967 , Narayanan1999 , Aminuddin2017 , Raju2017 , numerically Ranganayakulu1997 , Nellis2003 , Quintero2017 , Hasan2014 , and experimentally Gupta2000 , Yang2014 . Researchers have examined its effects on thermal performance under various conditions, emphasizing the importance of parameters such as thermal conductivity, plate thickness, and plate spacing Mitra2023 , Vera2018 . Optimization studies, such as those by Rogiers and Baelmans Rogiers2010 and Buckinx et al. Buckinx2013 , have further explored the effect of geometric parameters on the power density. These studies provide valuable insights but often assume ideal manufacturing conditions without considering AM-specific constraints.
In this paper, we investigate how material thermal properties and AM constraints influence the achievable compactness of counterflow plate heat exchangers. The study focuses on maximizing thermal power density under fixed pressure drop and effectiveness, while accounting for axial wall conduction effects inherent to compact designs. A simplified model of a parallel flat plate heat exchanger is used, with plate length and spacing as design variables and plate thickness treated as a constraint. The impact of AM-specific limitations on wall thickness and fouling risk is systematically evaluated across six different materials. Through this systematic study, we aim to contribute to the understanding of how material properties and their AM-specific limitations affect the compactness of cHEXs. It is shown that the AM-specific manufacturing limits can jeopardize the exploitation of the design freedom offered by this manufacturing technique for cHEXs.
@@ 2 Thermal model
The goal of this study is to minimize the size of a parallel-plate counterflow heat exchanger without compromising pressure drop and thermal effectiveness. For a given thermal power, this implies maximizing thermal power density. This section provides a comprehensive description of the analytical thermal model that is used to evaluate the power density, while accounting for axial heat conduction.
The present investigation is conducted on a simple parallel flat plate heat exchanger (PFPHE), as illustrated in Figure 1. The configuration comprises $2n+1$ flat, thin plates arranged in parallel, with each plate stacked on top of the preceding one. The plates have a thickness of $t$ and form $2n$ fluid channels in a repeating sequence, alternating between cold and hot flows. The channels are of identical spacing $D$, length $L$, and width $W$. Furthermore, this study assumes a perfectly balanced heat exchanger with equal mass flow rates, identical pressure drops, and uniform fluid properties on both the hot and cold sides.
@@
where $\varepsilon$ is the effectiveness, representing the ratio of the actual heat transfer rate to the maximum possible heat transfer rate Cengel2012 , $\dot{m}$ is the mass flow rate, $c_{p}$ is the fluid specific heat at constant pressure, and $\Delta T$ is the temperature difference between the hot and cold side inlets. Assuming a sufficiently large number of plates, the total height of the heat exchanger is approximated as $H=(2n)D+(2n+1)t\approx 2nD(1+t/D)$.
@@
where $f$ is the Fanning friction factor, Re is the Reynolds number based on hydraulic diameter $2D$, and $\mu$ is the dynamic viscosity. Upon substituting Eq. 3 into Eq 2, the mass flow rate $\dot{m}$ becomes
@@
By substituting Eq. 4 into Eq. 1 and applying algebraic manipulation followed by nondimensionalization, we derive an analytical expression for the nondimensional power density $Q^{\diamond}$:
$Q^{\diamond}\equiv\frac{\mathcal{D}\mu}{c_{p}\Delta T\rho\Delta P_{\text{ref}}}=\psi\frac{1}{f\textup{Re}}\frac{1}{1+t^{\ast}/D^{\ast}}\varepsilon\left(\frac{D^{\ast}}{L^{\ast}}\right)^{2}.$ (5)
Here, the nondimensional geometrical parameters are defined as
$L^{\ast}=\frac{L}{t_{\text{ref}}},\quad D^{\ast}=\frac{D}{t_{\text{ref}}},\quad t^{\ast}=\frac{t}{t_{\text{ref}}},$
where $t_{\text{ref}}$ is the reference plate thickness, and the factor $\psi$ is defined as the ratio of actual pressure drop $\Delta P$ to the reference pressure drop, i.e., $\psi=\Delta P/\Delta P_{\text{ref}}$. In this work, the reference values $t_{\text{ref}}$ and $\Delta P_{\text{ref}}$ correspond respectively to the plate thickness and pressure drop of the initial heat exchanger design, which is described in Table 1.
Kroeger Kroeger1967 derived an expression to calculate the effectiveness of a balanced counterflow heat exchanger, taking axial heat conduction in the wall into account:
@@
$\phi=\sqrt{\frac{\textup{M}\cdot\textup{NTU}}{1+\textup{M}\cdot\textup{NTU}}}\tanh{\frac{\textup{NTU}}{\sqrt{\textup{M}\cdot\textup{NTU}/(1+\textup{M}\cdot\textup{NTU})}}}.$ (7)
The number of transfer units NTU is a dimensionless parameter that combines the dimensional terms of the overall heat transfer coefficient $U$, heat transfer surface area $A$, mass flow rate $\dot{m}$, and specific heat $c_{p}$ Zhang2013 , i.e. $\textup{NTU}=\frac{UA}{\dot{m}c_{p}}$. Meanwhile, the axial conduction parameter $\textup{M}=\frac{UA_{ax}}{\dot{m}c_{p}}$ addresses the effect of heat conduction along the heat exchanger wall Kroeger1967 , Ranganayakulu1997 , with $A_{ax}$ is the cross-sectional area for axial heat conduction.
@@
where $h$ is the convective heat transfer coefficient, and $k_{w}$ is the thermal conductivity of the plate. The first term within the brackets accounts for the combined convective resistances of the hot and cold flow sides, while the second term addresses the lateral conduction resistance through the plate. By substituting the expression for the convective heat transfer coefficient $h=\frac{\textup{Nu}k}{2D}$ with Nu the Nusselt number and $k$ the fluid thermal conductivity, $UA$ is transformed into
$UA=2nk\frac{L}{D}\frac{\textup{Nu}}{4}W\left(1+\frac{\textup{Nu}}{4}\frac{k}{k_{w}}\frac{t}{D}\right)^{-1}.$ (9)
The axial plate conductance per unit length $UA_{ax}$, which is used in the calculation of the axial conduction parameter M, is expressed for the present geometry as
@@
By substituting Eqs. 4 and 9 into the definition of NTU, an analytical expression for the number of transfer units is obtained:
$\textup{NTU}=\frac{f\textup{Re}\textup{Nu}}{4}\left(1+\frac{\textup{Nu}}{4}\frac{k}{k_{w}}\frac{t}{D}\right)^{-1}\frac{k}{\rho c_{p}}\frac{\mu}{\Delta P}\frac{L^{2}}{D^{4}}.$ (11)
Likewise, by substituting Eqs. 4 and 10 into the definition of the axial conduction parameter M, we obtain:
$\textup{M}=f\textup{Re}\frac{k_{w}}{k}\frac{k}{\rho c_{p}}\frac{\mu}{\Delta P}\frac{t}{D^{3}}.$ (12)
Next, we express NTU and M in terms of the previously defined nondimensional parameters: $L^{\ast}=L/t_{\text{ref}},\;D^{\ast}=D/t_{\text{ref}},\;t^{\ast}=t/t_{\text{ref}}$, and $\psi=\Delta P/\Delta P_{\text{ref}}$. Additionally, we introduce a new dimensionless scaling group $\Pi=\frac{\Delta P_{\text{ref}}t_{\text{ref}}^{2}}{\alpha\mu}$, where $\alpha=\frac{k}{\rho c_{p}}$ is the fluid thermal diffusivity. The resulting forms of NTU and M are:
$\textup{NTU}=\frac{f\textup{Re}\textup{Nu}}{4}\left(1+\frac{\textup{Nu}}{4}\frac{k}{k_{w}}\frac{t^{\ast}}{D^{\ast}}\right)^{-1}\frac{1}{\psi\Pi}\frac{{L^{\ast}}^{2}}{{D^{\ast}}^{4}},$ (13)
$\textup{M}=f\textup{Re}\frac{k_{w}}{k}\frac{1}{\psi\Pi}\frac{t^{\ast}}{{D^{\ast}}^{3}}.$ (14)
These expressions allow for direct evaluation of the thermal effectiveness (via Eq. 6) and power density (via Eq. 5) as functions of the nondimensional geometrical and operating parameters, capturing the coupled effects of fluid flow, heat transfer, and axial conduction in a compact form. Together, they constitute an analytical model $Q^{\diamond}\left(D^{\ast},L^{\ast},t^{\ast}\right)$, as a function of the geometrical parameters $D^{\ast}$, $L^{\ast}$, and $t^{\ast}$.
This study considers heat exchangers operating in the laminar flow regime, which is characteristic of compact, low-Reynolds-number systems. The flow within each channel is assumed to be fully developed and two-dimensional (2D), confined between flat parallel plates. This fully developed flow assumption enables the use of constant values for the friction factor and Nusselt number, and is appropriate for channels with sufficiently long flow lengths relative to the hydraulic diameter. The 2D approximation is justified by the high width-to-spacing ratio that characterizes the geometry of the configuration studied here, which suppresses transverse velocity gradients and minimizes three-dimensional effects.
Under these conditions, the product of the Fanning friction factor and Reynolds number is constant, with $f\textup{Re}=24$ for laminar flow between infinite parallel plates Rohsenow1998 , Mikielewicz2014 , Lin2020 . Similarly, the Nusselt number is treated as a known constant. For a fully developed 2D planar channel flow subjected to constant surface heat flux on both walls, $\textup{Nu}=8.235$ Rohsenow1998 , Smith2014 . These values are used as fixed inputs in the present model and reflect typical conditions for idealized laminar channel flows. Additional 3D conjugate heat transfer (CHT) simulations have been conducted using ANSYS Fluent Ansys2024fluentr1guide to compare the analytical model predictions with high-fidelity numerical results. These results, provided in the supplementary material, show close agreement with the analytical predictions, with less than 0.52% relative deviation in thermal effectiveness (see Figure S1 for boundary conditions and Table S1 for simulation results in the supplementary material). This comparison supports the validity of the modeling assumptions adopted in this study.